OpenStudy (anonymous):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)? Enter your answer as a function beginning with f(x).

OpenStudy (cwrw238):

the simplest form is (x + 3)(x - 1) - this gives roots -3 and 1 now we need to know if vertex is at (-1,-8)

OpenStudy (anonymous):

can you give the f(x) of it?

OpenStudy (anonymous):

Format using this sample answer: f(x)=(3)(x^2-21x-19)

OpenStudy (cwrw238):

oh ok f(x) = (x^2 + 2x - 3)

OpenStudy (anonymous):

thank you.

OpenStudy (cwrw238):

to check if vertex is at (-1,-8) we write it in vertex form (x^2 + 2x - 3) = (x+ 1)^2 -4

OpenStudy (cwrw238):

- thats not the right answer- sorry but I'm a bit rusty ate these - ill check this out

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

right - to get vertex at (-1,8) the whole thing must be multiplied by 2 so its 2[(x + 1)^2 - 4] so correct answer is f(x) = (2)(x^2 + 2x - 3) sorry about that

OpenStudy (anonymous):

thank you that looks right!

OpenStudy (cwrw238):

yes - i've made sure using the math software wolfram alpha